The rate at which popcorn pops was measured at two different temperatures. How many times faster does the popcorn pop at 250.0 °C compared with that at 150.0 °C when the activation energy is 53.8 kJ/mol?
Answer #1

Activation energy = 53.8 kJ/mol

temperature T1 = 150 oC = 423 K

T2 = 250 oC = 523 K

ln (k2 / k1) = Ea / R [1 / T1 - 1/ T2]

ln (k2 / k1) = 53.8 / 8.314 x 10^-3 [1 / 423 - 1/523]

(k2 / k1) = 18.63

rate faster 18.6 times